# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

### v1

@@ Appendix A Supplementary proofs
$E_{1}$ $=2^{6-1/2r-r_{2}}dM^{2}_{\max}C_{2r},$
$E_{2}$ $=2^{4-2r}dM^{2}_{\max}C_{2r},$
$E_{3}$ $=2^{1+3s/2}dM^{2}_{\max}C_{s},$
$H(x)$ $=\underset{\phi\in\Phi}{\sup}\ \mathbb{E}\left[\left|\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right|^{x}\right]+1<\infty,\quad x\leq s$
[…]
where the latter one requires the additional assumption that the supremum of the ratio $a/b$ is bounded. Next, given for any $u>2$, for $|x-1|<1/2$ we arrive that for any $\tau\in(0,4]$:
$(x-1)^{\tau}\leq 2^{4-\min(u,4)}\left|x-1\right|^{\min(u,4)-(4-\tau)}.$ (33)
This proof follows [23], since the implicit constant is of interest, we still present it here for the sake of completeness.
Proof.
Consider the following event $\mathcal{A}$
where $S^{(a)}_{\ell-1}:=g^{(a)}_{\phi,M_{\ell-1}}(x)/Z(x,\phi)-1$, which is the sample mean of $M_{\ell-1}$ variable $\mathfrak{s}=g_{\phi}(x,\theta)/Z(x,\phi)-1$ with zero mean. Similarly, we define $S^{(b)}_{\ell-1}$ and $S_{\ell}:=g_{\phi,M_{\ell}}(x)/Z(x,\phi)-1$ and arrive at the following decomposition
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\right]=\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\mathbf{1}_{\mathcal{A}}\right]+\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\mathbf{1}_{\mathcal{A}^{c}}\right].$ (34)
We then need to develop a proper upper bound for this two terms. Denote the dimension of the $\phi$ as $d$, which is always finite, we leverage (32) to obtain
$\left\|\rho_{\phi,M_{\ell}}\right\|^{2}_{2}\leq\left\|\frac{\nabla_{\phi}g_{\phi}(x,\theta)}{g_{\phi}(x,\theta)}\right\|^{2}+\left\|\frac{\sum_{j=1}^{M_{\ell}}\nabla_{\phi}g_{\phi}(x,\theta_{j}^{\prime})}{\sum_{j=1}^{M_{\ell}}g_{\phi}(x,\theta_{j}^{\prime})}\right\|^{2}_{2}\leq 2dM_{\max}^{2},$
giving us a rough bound of the $\ell_{2}$ norm of $\rho_{\phi,M_{\ell}}$. That is
Then for $\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}$, we apply Jensen’s inequality to develop an upper bound
$\leq 3\left\|\rho_{\phi,M_{\ell}}\right\|_{2}^{2}+\frac{3}{4}\left\|\rho^{(a)}_{\phi,M_{\ell-1}}\right\|_{2}^{2}+\frac{3}{4}\left\|\rho^{(b)}_{\phi,M_{\ell-1}}\right\|_{2}^{2}\leq 9dM^{2}_{\max}.$
Then for the first term in (34) we have
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}^{2}\mathbf{1}_{\mathcal{A}}\right]\leq\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\mathbb{P}\left[\mathcal{A}\right]\leq 9dM^{2}_{\max}\mathbb{P}\left[\mathcal{A}\right].$ (35)
As for $\mathbb{P}[\mathcal{A}]$, applying Lemma 3.1 to have
$\mathbb{P}\left[\mathcal{A}\right]\leq\mathbb{P}\left[\left|S^{(a)}_{\ell-1}\right|>\frac{1}{2}\right]+\mathbb{P}\left[\left|S^{(b)}_{\ell-1}\right|>\frac{1}{2}\right]\leq\frac{2^{3s/2+1}C_{s}}{(M_{0}2^{\ell})^{s/2}}H(s),$
and it is proved that $\mathbb{E}[\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\mathbf{1}_{\mathcal{A}}]$ is of order $2^{-(s/2)\ell}$. For the second term $\mathbb{E}[\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\mathbf{1}_{\mathcal{A}^{c}}]$ in (34), utilizing the antithetic property to attain the following identity
$\Delta\rho_{\phi,\ell}$ $=\frac{1}{2}\left(\nabla g^{(a)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi)\right)\left(\frac{1}{g^{(a)}_{\phi,M_{\ell-1}}(x)}-\frac{1}{Z(x,\phi)}\right)$
$+\frac{1}{2}\left(\nabla g^{(b)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi)\right)\left(\frac{1}{g^{(b)}_{\phi,M_{\ell-1}}(x)}-\frac{1}{Z(x,\phi)}\right)$
$-\left(\nabla g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi)\right)\left(\frac{1}{g_{\phi,M_{\ell}}(x)}-\frac{1}{Z(x,\phi)}\right)$
$+\frac{1}{2}\frac{\nabla_{\phi}Z(x,\phi)}{g^{(a)}_{\phi,M_{\ell-1}}(x)}\left(\frac{g^{(a)}_{\phi,M_{\ell-1}}(x)}{Z(x,\phi)}-1\right)^{2}+\frac{1}{2}\frac{\nabla_{\phi}Z(x,\phi)}{g^{(b)}_{\phi,M_{\ell-1}}(x)}\left(\frac{g^{(b)}_{\phi,M_{\ell-1}}(x)}{Z(x,\phi)}-1\right)^{2}$
$-\frac{\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}\left(\frac{g_{\phi,M_{\ell}}(x)}{Z(x,\phi)}-1\right)^{2}.$ (36)
Now on event $\mathcal{A}^{c}$, we have $1/{g^{(a)}_{\phi,M_{\ell-1}}(x)}\leq 2/{Z(x,\phi)},1/{g^{(b)}_{\phi,M_{\ell-1}}(x)}\leq 2/{Z(x,\phi)}$. With simple algebra, we have $1/{g_{\phi,M_{\ell}}(x)}\leq 2/{Z(x,\phi)}$. Also we have $|S_{\ell}|<1/2$ on $\mathcal{A}^{c}$. Then an upper bound of the normalizing constant estimators on event $\mathcal{A}^{c}$ are given. We denote $F^{(a)}_{\ell-1}:=(\nabla g^{(a)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi))/Z(x,\phi),\tilde{F}^{(a)}_{\ell-1}:=(\nabla g^{(a)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi))/g^{(a)}_{\phi,M_{\ell-1}}$. Where the former one is also a mean of $M_{\ell-1}$ variable $\mathfrak{f}:=\nabla_{\phi}g_{\phi}(x,\theta)-\nabla_{\phi}Z(x,\phi)/Z(x,\phi)$ with zero mean. Similarly, we define $F^{(b)}_{\ell-1},\tilde{F}^{(b)}_{\ell-1}$ and $F_{\ell}:=(\nabla g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi))/Z(x,\phi),\tilde{F}_{\ell}:=(\nabla g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi))/g_{\phi,M_{\ell}}(x)$. Directly applying Jensen’s inequality on (36) to have
@@
$+2\left\|\frac{\nabla_{\phi}Z(x,\phi)}{g^{(a)}_{\phi,M_{\ell-1}}(x)}\right\|_{2}^{2}(S^{(a)}_{\ell-1})^{4}+2\left\|\frac{\nabla_{\phi}Z(x,\phi)}{g^{(b)}_{\phi,M_{\ell-1}}(x)}\right\|_{2}^{2}(S^{(b)}_{\ell-1})^{4}$
$+4\left\|\frac{\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}\right\|_{2}^{2}(S_{\ell})^{4}.$ (37)
With the upper bound derived before, we arrive at the following:
@@
$+8\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2}^{2}\left((S^{(a)}_{\ell-1})^{4}+(S^{(b)}_{\ell-1})^{4}+2(S_{\ell})^{4}\right),$ (38)
where each term is some power of a mean of $\mathfrak{f}$ or $\mathfrak{s}$ enabling one to use Lemma 3.1 to derive their upper bound. To obtain the order of $\mathbb{E}[\|\Delta\rho_{\phi,l}\|_{2}^{2}\mathbf{1}_{\mathcal{A}^{c}}]$, we begin by applying (32) to yield
$\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2}^{2}=\left\|\frac{\mathbb{E}[\nabla_{\phi}g_{\phi}(x,\theta)]}{\mathbb{E}[g_{\phi}(x,\theta)]}\right\|_{2}^{2}\leq\underset{x,\theta,\phi}{\mathrm{sup}}\left\|\frac{\nabla_{\phi}g(x,\theta)}{g_{\phi}(x,\theta)}\right\|_{2}^{2},$ (39)
then it would be sufficient to derive an upper bound for the expectation of the $\|F_{\ell}\|_{2}^{2}(S_{\ell})^{2}$ and $(S_{\ell})^{4}$ in (38) on event $\mathcal{A}^{c}$ due to the symmetry. We firstly start from $\|F_{\ell}\|_{2}^{2}(S_{\ell})^{2}$, applying (33) to obtain
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\left(S_{\ell}\right)^{2}\textbf{1}_{\mathcal{A}^{c}}\right]$ $\leq\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\times 2^{4-2r}\left|S_{\ell}\right|^{{2r-2}}\right]$
$\leq\left(\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]\right)^{1/r_{2}}2^{4-2r}\left(\mathbb{E}\left[\left|S_{\ell}\right|^{{2r}}\right]\right)^{1-1/r_{2}}.$ (40)
@@
To obtain the desired upper bound, we handle with these two moments separately. For $\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]$, with Jensen’s inequality and Lemma 3.1, we have
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]\leq\frac{C_{2r}}{M_{\ell}^{r_{2}}}\mathbb{E}\left[\|\mathfrak{f}\|^{2r}_{2}\right],$
then for $\mathbb{E}\left[\|\mathfrak{f}\|^{2r}_{2}\right]$ on $\mathcal{A}^{c}$, applying Jensen’s inequality
$\mathbb{E}\left[\|\mathfrak{f}\|^{2r}_{2}\right]$ $=\mathbb{E}\left[\left\|\frac{\nabla_{\phi}g_{\phi}(x,\theta)-\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|^{2r}_{2}\right]$
$\leq d^{r_{2}-1}2^{2r-1}\mathbb{E}\left[\left\|\frac{\nabla_{\phi}g_{\phi}(x,\theta)}{g_{\phi}(x,\theta)}\cdot\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right\|_{2r}^{2r}+\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2r}^{2r}\right].$
Then an upper bound for $\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]$ in (40) is presented as
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r}\right]\leq\frac{2^{2r-1}d^{r_{2}}M^{2r}_{\max}C_{2r}}{M_{\ell}^{r_{2}}}H(2r).$
For the upper bound of $\mathbb{E}\left[\left|S_{\ell}\right|^{{2r}}\right]$ in (40), we directly apply Lemma 3.1 to have
$\mathbb{E}\left[\left|S_{\ell}\right|^{{2r}}\right]\leq\frac{C_{2r}}{M_{\ell}^{r_{2}}}H(2r),$
directly plugging them into (40) to obtain the desired upper bound
$\mathrm{RHS}=\left(2^{2r-1}d^{r_{2}}M^{2r}_{\max}C_{2r}\right)^{1/r_{2}}2^{4-2r}C^{1-1/{r_{2}}}_{2r}\frac{H(2r)}{(M_{0}2^{\ell})^{r_{2}}}=\frac{E_{1}H_{2r}}{(M_{0}2^{\ell})^{r_{2}}}.$
@@
With similar treatment for $\mathbb{E}[\|F^{(a)}_{\ell-1}\|_{2}^{2}(S^{(a)}_{\ell-1})^{2}1_{\mathcal{A}^{c}}]$ and $\mathbb{E}[\|F^{(b)}_{\ell-1}\|_{2}^{2}(S^{(b)}_{\ell-1})^{2}1_{\mathcal{A}^{c}}]$, we then have a uniform upper bound as $E_{1}H_{2r}2^{r_{2}}/(M_{0}2^{\ell})^{r_{2}}$.
As for the $\mathbb{E}[(S_{\ell})^{4}1_{\mathcal{A}^{c}}]$, the sixth term in (38), note that
$\mathbb{E}[(S_{\ell})^{4}1_{\mathcal{A}^{c}}]\leq 2^{4-2r}\mathbb{E}\left[\left|S_{\ell}\right|^{2r}\right]\leq 2^{4-2r}\frac{C_{2r}}{M_{\ell}^{r_{2}}}H(2r)$

### v2

@@ 6 Concluding remarks
$\tilde{F}^{(i)}_{\ell-1}$ $:=\left(\nabla_{\phi}g^{(i)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi)\right)/g^{(i)}_{\phi,M_{\ell-1}}(x),\ i\in\{a,b\},$
$\tilde{F}_{\ell}$ $:=\left(\nabla_{\phi}g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi)\right)/g_{\phi,M_{\ell}}(x).$
We denote
$F^{(i)}_{\ell-1}$ $:=\left(\nabla g^{(i)}_{\phi,M_{\ell-1}}(x)-\nabla_{\phi}Z(x,\phi)\right)/Z(x,\phi),\ i\in\{a,b\},$
$F_{\ell}$ $:=\left(\nabla g_{\phi,M_{\ell}}(x)-\nabla_{\phi}Z(x,\phi)\right)/Z(x,\phi),$
which are sample means of the random variable
$\mathfrak{f}:=\left(\nabla_{\phi}g_{\phi}(x,\theta)-\nabla_{\phi}Z(x,\phi)\right)/Z(x,\phi)$
with zero mean. Directly applying Jensen’s inequality on (41) gives
@@
$+2\left\|\frac{\nabla_{\phi}Z(x,\phi)}{g^{(a)}_{\phi,M_{\ell-1}}(x)}\right\|_{2}^{2}(S^{(a)}_{\ell-1})^{4}+2\left\|\frac{\nabla_{\phi}Z(x,\phi)}{g^{(b)}_{\phi,M_{\ell-1}}(x)}\right\|_{2}^{2}(S^{(b)}_{\ell-1})^{4}+4\left\|\frac{\nabla_{\phi}Z(x,\phi)}{g_{\phi,M_{\ell}}(x)}\right\|_{2}^{2}(S_{\ell})^{4}.$
Now on event $\mathcal{A}^{c}$, we have $1/{g^{(a)}_{\phi,M_{\ell-1}}(x)}\leq 2/{Z(x,\phi)},1/{g^{(b)}_{\phi,M_{\ell-1}}(x)}\leq 2/{Z(x,\phi)}$. With simple algebra, we have $1/{g_{\phi,M_{\ell}}(x)}\leq 2/{Z(x,\phi)}$. Also we have $|S_{\ell}|\leq 1/2$ on $\mathcal{A}^{c}$. With the upper bound derived before, on event $\mathcal{A}^{c}$, we arrive at the following:
@@
$\quad+8\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{2}^{2}\left((S^{(a)}_{\ell-1})^{4}+(S^{(b)}_{\ell-1})^{4}+2(S_{\ell})^{4}\right).$ (42)
Applying Jensen’s inequality for any $u>0$ yields
$\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{u}^{u}=\frac{\left\|\mathbb{E}[\nabla_{\phi}g_{\phi}(x,\theta)|x]\right\|_{u}^{u}}{Z(x,\phi)^{u}}\leq\frac{dM_{\max}^{u}\mathbb{E}_{[}g_{\phi}(x,\theta)|x]^{u}}{Z(x,\phi)^{u}}=dM_{\max}^{u}.$ (43)
By setting $u=2$, it is sufficient to derive an upper bound for the expectation of the $\|F_{\ell}\|_{2}^{2}(S_{\ell})^{2}$ and $(S_{\ell})^{4}$ in (42) on event $\mathcal{A}^{c}$, since directly setting $\ell$ to $\ell-1$ gives the upper bound for the rest of the terms. By (35) and (43), we have
$\mathbb{E}\left[\|\mathfrak{f}\|^{u}_{u}\right]$ $=\mathbb{E}\left[\left\|\frac{\nabla_{\phi}g_{\phi}(x,\theta)-\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|^{u}_{u}\right]$
$\leq 2^{u-1}\mathbb{E}\left[\left\|\frac{\nabla_{\phi}g_{\phi}(x,\theta)}{Z(x,\phi)}\right\|_{u}^{u}+\left\|\frac{\nabla_{\phi}Z(x,\phi)}{Z(x,\phi)}\right\|_{u}^{u}\right]$
$\leq 2^{u-1}dM_{\max}^{u}\left(\mathbb{E}\left[\left(\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right)^{u}\right]+1\right)$
$\leq d2^{u-1}M_{\max}^{u}(H(u)+1).$ (44)
If $|x-1|\leq 1/2$, we then have for any $\tau\in(0,4]$,
$|x-1|^{\tau}\leq 2^{\min(s,4)-4}\left|x-1\right|^{\min(s,4)-(4-\tau)}=2^{2r_{2}-4}\left|x-1\right|^{2r_{2}-(4-\tau)},$ (45)
where $r_{2}=\min(s,4)/2$. Applying (45) with $\tau=2$ and Hölder inequality obtains
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\left(S_{\ell}\right)^{2}\textbf{1}_{\mathcal{A}^{c}}\right]$ $\leq\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\times 2^{2r_{2}-4}\left|S_{\ell}\right|^{{2r_{2}-2}}\right]$
$\leq 2^{2r_{2}-4}\left(\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2r_{2}}\right]\right)^{1/r_{2}}\left(\mathbb{E}\left[\left|S_{\ell}\right|^{{2r_{2}}}\right]\right)^{1-1/r_{2}}.$ (46)
@@
To obtain the desired upper bound, we handle these two moments separately. Since $\mathbb{E}\left[\|\mathfrak{f}\|^{u}_{u}\right]$ is bounded by (A), applying Jensen’s inequality with Lemma A.1 gives
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{u}\right]\leq\frac{d^{u/2-1}C_{u}}{M_{\ell}^{u/2}}\mathbb{E}\left[\|\mathfrak{f}\|^{u}_{u}\right]\leq\frac{d^{u/2}2^{u-1}C_{u}M_{\max}^{u}(H(u)+1)}{M_{\ell}^{u/2}}.$ (47)
We directly apply Lemma A.1 to have
$\mathbb{E}\left[\left|S_{\ell}\right|^{{u}}\right]\leq\frac{C_{u}}{M_{\ell}^{u/2}}\mathbb{E}[|g_{\phi}(x,\theta)/Z(x,\phi)-1|^{u}]\leq\frac{C_{u}}{M_{\ell}^{u/2}}(H(u)+1).$ (48)
Using (46), (47) and (48) gives
$\mathbb{E}\left[\left\|F_{\ell}\right\|_{2}^{2}\left(S_{\ell}\right)^{2}\textbf{1}_{\mathcal{A}^{c}}\right]=\mathcal{O}(M_{\ell}^{-r_{2}}).$
@@
With similar treatment for $\mathbb{E}[\|F^{(a)}_{\ell-1}\|_{2}^{2}(S^{(a)}_{\ell-1})^{2}\mathbf{1}_{\mathcal{A}^{c}}]$ and $\mathbb{E}[\|F^{(b)}_{\ell-1}\|_{2}^{2}(S^{(b)}_{\ell-1})^{2}\mathbf{1}_{\mathcal{A}^{c}}]$, we then have an upper bound of $\mathcal{O}(M_{\ell}^{-r_{2}})$ for them.
Applying (45) with $\tau=4$, we have
$\mathbb{E}[(S_{\ell})^{4}\mathbf{1}_{\mathcal{A}^{c}}]\leq 2^{2r_{2}-4}\mathbb{E}\left[\left|S_{\ell}\right|^{2r_{2}}\right]\leq 2^{2r_{2}-4}\frac{C_{2r_{2}}}{M_{\ell}^{r_{2}}}(H(2r_{2})+1).$